Site Conditions:
  • No smoke is observable
  • Odor still slightly noticeable
  • Erosion controls appear to be effective with minimal erosion observed
  • Grading and filling operations on the western and northern portions of the site are currently underway
  • Weather permitting, approximately 150 loads of fill dirt will be brought to the site today
  • Stone continues to be placed around the site to reduce runoff and erosion